OpenStudy (anonymous):

Jason ran for 1.5 hours with a 2 mph wind at his back. It took him 2.5 hours to get home with the same wind in his face. How would Jason have run if not for the wind?

OpenStudy (shubhamsrg):

sorry did above one wrongly.. the eqns will be v + vw = d/1.5 and v - vw = d/2.5 now seems correct..

OpenStudy (shubhamsrg):

v= his speed vw=2(wind speed) d= distance..

OpenStudy (anonymous):

So would i use something like elimination or substitution on them or something ? @shubhamsrg

OpenStudy (shubhamsrg):

yep..just solve for v,,you may use any legal method..

OpenStudy (anonymous):

I dont get what Im supposed to do. it says 2mph with wind at his back so is that the speed or wtf ..

OpenStudy (anonymous):

you can think of them as forces i guess

OpenStudy (anonymous):

however i'm not sure if those are correct

OpenStudy (anonymous):

velocityor speed has a quantity of distance per time and the wind also has it so you have